# NGO Activism:

### v7

@@ Appendix B Omitted Figures
Figure B1: Persistence of the AGM month over time
Note: This figure plots the percentage of AGMs happening in each month of the year (y-axis), given the month of the previous year’s AGM (x-axis). A larger and darker circle indicates a greater proportion of meetings. Plot based on 38,471 meetings of 5,745 US companies, from 2010 to 2020.
Figure B2: Firms’ negative media coverage increases if NGOs campaign during AGMs
Note: The figure plots the coefficients $\beta_{\tau}$ from the following event study regression where we re-calendarize time around an AGM date so that the AGM month is month 0 in each year for each firm:
$News_{n,w,y}$ $=\sumop\displaylimits_{\begin{subarray}{c}\tau=-5\\ \tau\neq-4\end{subarray}}^{5}\!\left(\beta_{\tau}\cdot 1\!\left\{M_{n,w,y}=\tau\!\right\}\cdot NGO\,Campaign_{n,w,y}\!\right)$
$\qquad+X_{n,w,y}\gamma^{\prime}+\alpha_{n,y}+\alpha_{w,y,s(n)}+u_{n,w,y},$
where $News_{n,w,y}$ is an indicator for whether firm $n$ in week $w$ of year $y$ experiences at least one negative E&S news event. The set of dummies $1\{M_{n,w,y}=\tau\}$ indicate the distance in months from the closest AGM. The function $M_{n,w,y}$ assigns week $w$ to each month; the indicator is 1 if the week falls in month $\tau$, where a month is made up of 4 weeks, and zero otherwise. We interact these event-study indicators, with the variable $NGO\,Campaign_{n,w,y}$, an indicator variable equal to 1 if firm n experienced an NGO campaign in week $w$ of year $y$, and 0 otherwise. The vector $X_{n,w,y}$ includes the direct effect of these variables. The reference month is $\tau=-4$ following SEC rules on shareholder proposals. Note that, because of the stability of AGM dates across years within firms (Appendix Figure B1), the distribution of months fixed effects is unaffected by the re-calendarization, so that months are not double counted across years. Standard errors are clustered by firm.
Figure B3: NGO campaigns and shareholder proposals
@@
Note: This figure plots the estimated coefficients of the regression of a dummy indicating the occurrence of a shareholder proposal on month $t$ at firm $f$, relative to an NGO campaign targeting firm $f$ on the same cause, $k$ months ahead. The regression also includes campaign and month-by-year-by-industry-by-topic fixed effects. The regression is based on 1,227 NGO campaigns. Standard errors are clustered at the firm level.
Figure B4: NGO campaign timing and consumer backlash
(a) NGO Campaigns on AGM dates
(b) NGO Campaigns on Other Days
Notes: The figure plots the estimated ${\beta}_{\tau}$ from (3). Panel (a) focuses on NGO campaigns on AGM dates, while Panel (b) focuses on campaigns on dates other than AGM dates. The reference period is -1 quarter, consistent with the other analyses. The number of observations is 239,818 in Panel (a) and 239,925 in Panel (b). Standard errors are clustered at the firm level.
Figure B5: Campaign Timing and E&S Score Improvement
Note: This figure plots the estimated coefficients of the following regression:
$_{5}\text{E\&S Score}_{n,q}=\sumop\displaylimits_{l=-6}^{5}\text{Campaign}_{n,q,l}+X_{n,q}\gamma^{\prime}+\alpha_{n}+\alpha_{s(n),q}+u_{n,q},$ (B1)
where $_{5}\text{E\&S Score}_{i,q}$ is the change in firm $n$’s average of its Environmental and Social scores between quarter q1q-1 and quarter $q+4$, and $\sumop\displaylimits_{l=-6}^{6}\text{Campaign}_{n,q,l}$ is a set of dummy variables indicating the occurrence of an NGO campaign against firm $n$ in quarter $q$, being $l$ months away from the closest AGM. The regression also controls ($X_{n,q}$) for firm size (logged total assets), lagged average of E and S scores, firm fixed effects ($\alpha_{n}$), and industry-by-quarter fixed effects ($\alpha_{s(n),q}$). ESG scores are from MSCI, NGO campaigns data from Sigwatch, firm characteristics from Compustat, and AGM dates from ISS. The regression has 47,082 observations. Standard errors are clustered at the firm level.
Unlike an event study, equation (B1) is a distributed leads-and-lags regression. We omit a reference period, so the coefficients are interpreted relative to firms that did not experience a campaign. This choice allows us to directly compare the effects of early campaigns with those launched on the AGM date relative to firms without campaigns. Results remain similar if we instead set month –4 as the reference period, as in the other analyses.
@@
Note: This figure plots the estimated coefficients from regressing the target firm’s market cap (in deciles to control for the fact that the tails of the market cap distribution have thickened in recent years) on dummies indicating the number of campaigns the focal NGO had launched in previous years as in (5). Standard errors are clustered at the firm level. The number of observations is 1,338.
@@
(a) Target firm’s E score for Environmental campaigns
(b) Target firm’s S score for Social campaigns
Note: This figure plots the estimated coefficients from regressing either a target firm’s environmental score (top panel) or its social score (bottom panel) on dummies indicating either the number of environmental (top panel) or social (bottom panel) campaigns the focal NGO had launched in previous years as in (5). Standard errors are clustered at the firm level. The number of observations is 636 and 698, respectively.
Figure B10: Media coverage of NGO-target firms around their AGM
Note: To test if media exposure of NGO campaigns is highest during the AGM week, we aggregate negative E&S news by firm ($n$) and week-by-year ($w,y$) and perform the event study:
$\text{News}_{n,w,y}$ $=\sumop\displaylimits_{\begin{subarray}{c}\tau=-5\\ \tau\neq-4\end{subarray}}^{5}\left(\beta_{\tau}\cdot\mathbf{1}_{\{M_{n,w,y}=\tau\}}\cdot\text{NGO Campaign}_{n,w,y}\right)$
$\qquad+X_{n,w,y}\gamma^{\prime}+\alpha_{n,y}+\alpha_{w,y,s(n)}+u_{n,w,y},$
[…]
@@ Appendix C Robustness of Google Trends Results
This section discusses the robustness of the results presented in Figure 2 to alternative specifications. In particular, we replicate the analysis using (i) a staggered difference-in-differences estimator (Sun and Abraham, 2021), which accounts for heterogeneous campaign timing across NGOs, and (ii) a distributed lags specification, which controls for the possibility of overlapping campaigns within the same year.
@@ C.1 Staggered Difference-in-Differences Specification
We first replicate Figure 2, applying the staggered difference-in-differences approach to account for heterogeneous campaign timing across NGOs (Sun and Abraham, 2021).
Estimated coefficients $\beta_{\tau}^{AGM}$ are reported in Figure C1. The results are consistent with our baseline findings: AGM-timed campaigns (blue dots) generate sizeable visibility gains for early campaigns, raising Google search intensity by about 58%. By contrast, NGOs see a more modest increase (33%) in search intensity for later campaigns, and only in the month of the campaign.
@@
Note: The Figure plots the estimated $\beta^{AGM}$ coefficients from (1) using a staggered difference-in-differences estimator (Sun and Abraham, 2021). The dependent variable is the monthly Google Trends score for the NGO. Blue dots are estimated by restricting the sample to an NGO’s first three campaigns, while red triangles consider later campaigns. Standard errors are clustered at the campaign level.

### v8

@@ Appendix B Omitted Figures
Figure B1: How does target characteristics affect the decision to campaign on AGM dates?
Note: The figure reports the estimated coefficients of the regression of an indicator that an NGO campaign was launched on the target firm’s AGM date on the displayed (standardized) variables, and year fixed effects. Error bars shows 90% (shaded) and 95% confidence intervals using standard errors that are clustered by firm.
Figure B2: Increased NGO visibility around an AGM campaign – All sample
(a) Campaign timing and campaigning NGO’s Google Trends scores
(b) Campaign timing and number of target firms’ negative news (standardized)
Note: Each panel plots the estimated $\beta^{AGM}$ coefficients from (10). In Panel (a), the dependent variable is the monthly Google Trends score for the NGO. In Panel (b), the dependent variable is the (normalized) number of monthly negative media coverage received by a target firm. Error bars shows 95% confidence intervals using standard errors that are clustered by NGO.
Figure B3: NGO campaigns and shareholder proposals – All Sample
(a) All NGOs
@@
Notes: Note: This figure plots the estimated coefficients of the regression of a dummy indicating the occurrence of a shareholder proposal on month $t$ at firm $f$, relative to an NGO campaign targeting firm $f$ on the same cause, $k$ months ahead. The regression also includes campaign and month-by-year-by-industry-by-topic fixed effects. Error bars shows 95% confidence intervals using standard errors that are clustered by firm.
Figure B4: NGO campaign timing and donations – All sample
(a) Donations
(b) Operating Margin
Notes: The figure plots the estimated ${\beta}_{\tau}$ from (12). Panel (a) focuses on donations over (lagged) assets, while Panel (b) focuses on donations minus expenditures over (lagged) assets. Error bars shows 95% confidence intervals using standard errors that are clustered by NGO.
Figure B5: Campaign Timing and Changes in E&S Scores – All Sample
Note: The figure plots the estimated coefficients $\beta_{l}$ from equation (15). The dependent variable is the yearly change in a firm’s average E&S scores. Shaded areas show 95% confidence intervals based on firm-level clustered standard errors. Data combine MSCI ESG scores, NGO campaigns from Sigwatch, firm characteristics from Compustat, and AGM dates from ISS.
@@
Note: This figure plots the estimated coefficients from regressing the target firm’s market cap (in deciles to control for the fact that the tails of the market cap distribution have thickened in recent years) on dummies indicating the number of campaigns the focal NGO had launched in previous years as in (16). Standard errors are clustered at the firm level.
@@
(a) Target firm’s E score
(b) Target firm’s S score
Note: This figure plots the estimated coefficients from regressing either a target firm’s environmental score (top panel) or its social score (bottom panel) on dummies indicating either the number of environmental (top panel) or social (bottom panel) campaigns the focal NGO had launched in previous years as in (16). Standard errors are clustered at the firm level.
@@ Appendix C Robustness of Google Trends Results
This section discusses the robustness of the results presented in Figure 3 to alternative specifications. In particular, we replicate the analysis using (i) a staggered difference-in-differences estimator (sun2021estimating), which accounts for heterogeneous campaign timing across NGOs, and (ii) a distributed lags specification, which controls for the possibility of overlapping campaigns within the same year.
@@ C.1 Staggered Difference-in-Differences Specification
We first replicate Figure 3, applying the staggered difference-in-differences approach to account for heterogeneous campaign timing across NGOs (sun2021estimating).
Estimated coefficients $\beta_{\tau}^{AGM}$ are reported in Figure C1. The results are consistent with our baseline findings: AGM-timed campaigns (blue dots) generate sizeable visibility gains for young NGOs (first four campaigns), raising Google search intensity by about 58%. By contrast, NGOs see a more modest increase (33%) in search intensity for later campaigns, and only in the month of the campaign.
@@
Note: The Figure plots the estimated $\beta^{AGM}$ coefficients from (10) using a staggered difference-in-differences estimator (sun2021estimating). The dependent variable is the monthly Google Trends score for the NGO. Blue dots are estimated by restricting the sample to an NGO’s first three campaigns, while red triangles consider later campaigns. Standard errors are clustered at the campaign level.